OpenStudy (anonymous):

find the 15th term in the geometric sequence below. 20,10,5,2.5,... A: -(25/4096) B: -(5/4096) C: 5/4096 D: 25/4096

Parth (parthkohli):

\( \color{Black}{\Rightarrow \large a_1 \times r^{n - 1}}\) If you have a teacher, then use this crappy formula.

OpenStudy (anonymous):

First term=20 R=0.5 Tn=ar^n-1 where a=first term , r=ratio ,n=number of the term

Parth (parthkohli):

\( \color{Black}{\Rightarrow 20 \times 0.5^{14}}\)